### v1

# Reading Subtext: Evaluating Large Language Models on Short Story Summarization with Writers

Abstract
We evaluate recent Large language Models (LLMs) on the challenging task of summarizing short stories, which can be lengthy, and include nuanced subtext or scrambled timelines. Importantly, we work directly with authors to ensure that the stories have not been shared online (and therefore are unseen by the models), and to obtain informed evaluations of summary quality using judgments from the authors themselves. Through quantitative and qualitative analysis grounded in narrative theory, we compare GPT-4, Claude-2.1, and LLama-2-70B. We find that all three models make faithfulness mistakes in over 50% of summaries and struggle to interpret difficult subtext. However, at their best, the models can provide thoughtful thematic analysis of stories. We additionally demonstrate that LLM judgments of summary quality do not match the feedback from the writers.
1 Introduction
Narrative is a fundamental part of how we communicate and make sense of our experiences. As Herman (2009) describes, “Narrative roots itself in the lived, felt experience of human or human-like agents interacting in an ongoing way with their surrounding environment…" Understanding narrative is, therefore, a necessary skill large language models (LLMs) need in order to engage with the subtleties of human experience and communication. We test how well LLMs understand the subtleties of narrative through the task of narrative summarization, with a focus on interpreting themes and meaning in addition to capturing plot.
For our narrative form, we use fiction short stories as they present several interesting challenges. Fiction writing does not follow a clear intro-body-conclusion format and instead may follow a nonlinear timeline, hint at things only abstractly, or deliberately mislead the reader. Fiction may use multiple language varieties and creative language. For example, consider this quote from Toni Morrison’s Beloved: “The pieces I am, she gather them and give them back to me in all the right order" Morrison (2004). This sentence uses complex metaphor and African American Language Deas et al. (2023); Grieser (2022) from the late 1800’s to express a beautiful relationship between two characters. Finally, fiction short stories can be longer than an LLM’s context window. Here, we consider stories up to 10,000 tokens long.
Evaluations of narrative summarization on long documents have been scarce due to several key challenges: 1) Narrative text is generally either in the public domain (and therefore likely in LLM training data) or under copyright, and 2) Holistic summary evaluation has been prohibitively difficult due to a lack of reliable automatic metrics Fabbri et al. (2021a); Chang et al. (2023) and complications with human evaluation. For example, it can take someone over an hour to thoroughly read and evaluate just one story and summary, which quickly becomes expensive.
Chang et al. (2023) make progress on these issues by purchasing recent books (which are less likely to be in models’ training data but may still be discussed in the training data), and developing an LLM-based metric for evaluating summary coherence. We take this a step further by working directly with experienced creative writers, and thus are able to: 1) use stories that are not discussed or present in training data, 2) expand beyond evaluation of coherence to attributes focused on narrative understanding like faithfulness and thematic analysis, and 3) use human rather than model judgments.
We evaluate three LLMS – GPT-4, Claude-2.1, and LLama-2-70b – on 25 short stories. We ask authors to evaluate the summaries of their own unpublished stories [^1] We release our code, the writer evaluation responses, and the story/summary errors here: https://github.com/melaniesubbiah/reading-subtext. since they are experts in what they have written, focusing on coherence, faithfulness, coverage of important details, and useful interpretation in the summaries. We then present an analysis of the quantitative and qualitative feedback. Our contributions include:
  1. 1.
    A characterization of LLM narrative understanding based on narrative theory Mandler and Johnson (1977); Piper et al. (2021), resulting in directions for future research.
  2. 2.
    A methodology for evaluating LLM summaries when the lines between test and training data have blurred, as well as a demonstration that LLM judgments cannot replace skilled human evaluators for this task.
  3. 3.
    Informed ratings of faithfulness and the interpretive quality of LLM summaries using experienced creative writers and unseen short stories.
Lastly, this work demonstrates the mutual benefit of working directly with communities who have valuable data as an increasing amount of online content is consumed or generated by models.
2 Related Work
Narrative Summarization. Interest in long narrative summarization has steadily grown in the last several years. Ladhak et al. (2020) first introduced the task of summarizing chapters from novels, which was then expanded into the full BookSum dataset Kryściński et al. (2021) and used in early studies of RLHF Wu et al. (2021). Most similar to our work is SQuALITY Wang et al. (2022), which also focuses on short stories. However, their stories are sourced from Project Gutenberg which is likely memorized by LLMs. Across other areas of narrative understanding, recent datasets have been proposed in screenplays Chen et al. (2021), poetry Mahbub et al. (2023), and theory of mind evaluation Xu et al. (2024).
Summarization Evaluation. Evaluating summary quality in any of these areas is a challenge. Work like Fabbri et al. (2021a) has shown the flaws in many of the traditional automatic metrics, prompting a move toward model-based reference-free methods and fine-grained span analysis. LongEval Krishna et al. (2023), QAFactEval Fabbri et al. (2021b), FALTE Goyal et al. (2022a), and FActScore Min et al. (2023) have furthered methods in faithfulness, while work like SNaC Goyal et al. (2022b) and BooookScore Chang et al. (2023) have focused on coherence. Chang et al. (2023) is most similar to our work, but they focus on a GPT prompting strategy for evaluation, evaluate on recently published books, and only evaluate summary coherence. Several studies have benchmarked LLMs on summarization tasks Tang et al. (2024); Zhang et al. (2023a); Tang et al. (2022); Jahan et al. (2024); Liu et al. (2023); Pu et al. (2023), including ones Zhang et al. (2023b); Goyal et al. (2023) that have shown performance is saturated on the commonly used CNN/DM news summarization dataset Hermann et al. (2015).
Studies with Writers. There have been some studies that collaborate with writers to address the evaluation problem. However, these all focus on narrative generation or collaborative writing and most use amateur writers or crowdworkers rather than skilled professionals Zhong et al. (2023); Yuan et al. (2022); Begus (2023); Chakrabarty et al. (2022); Padmakumar and He (2023); Yeh et al. (2024). Most relevant to our work are these studies that also use professional writers but focus on narrative generation instead of summarization:
Chakrabarty et al. (2023a) – 10 writers, 12 stories
Chakrabarty et al. (2023b) – 17 writers, 30 stories
Ippolito et al. (2022) – 13 writers, 13 stories
Huang et al. (2023) – 8 writers, 8 stories
These studies use similar numbers of writers and stories to us given the challenges/cost involved.
3 Writers and Data
Given the problems with using stories that may have contents or analysis available online, we want to find skilled writers with unpublished stories they are willing to share. We therefore recruit writers through MFA listservs, posts on X, and direct emails. Once writers express interest, we screen for skill by writer education level and/or portfolio. Finally, we obtain their informed consent by sharing an infosheet on the study, which explains the task of sharing their stories and evaluating summaries, and includes how their work will be used as outlined in IRB protocol AAAU8875 (see full infosheet in Appendix A). We ask two questions to determine the background of the writers: 1) Do you have an undergraduate or MFA writing degree?, and 2) Have you previously published any of your writing? To protect writers’ anonymity, we use anonymous IDs to collect and store all the data. In line with Chakrabarty et al. (2023a) and Ippolito et al. (2022), who use 8 and 13 writers respectively, our recruited group consists of 9 skilled writers – 8/9 possess a writing degree, and 7/9 have previously published writing.
Each writer is given the opportunity to submit short stories that they have written and not published anywhere online. We limit the writers to five stories at most, so one writer does not dominate the data. Four writers choose to submit five stories, and five submit one story, resulting in a dataset of 25 stories. We compensate the writers for their submitted stories and evaluations. Table 1 shows some summary statistics for this data. None of the stories exceed 9900 tokens in length, and for the purpose of analysis, we consider performance across three length buckets: short (<3300 tokens), medium (3300-6600 tokens), and long (6600-9900 tokens). Token count is computed as the average of the GPT-4 tokenizer [^2] https://cookbook.openai.com/examples/how_to_count_tokens_with_tiktoken count and the Llama2 tokenizer [^3] https://huggingface.co/docs/transformers/model_doc/llama2#transformers.LlamaTokenizer count as Claude does not have a publicly available tokenizer.
4 Summary Generation
Figure 1: The two different methods used for summarization and the associated prompts for the models. GPT-4 and Claude have sufficient input context to summarize a whole story, whereas Llama has to use a chunk-then-summarize approach for longer stories.
Length Stories Summary Avg. Len.
Bucket # Avg. Len. GPT4 Claude Llama
Short 10 1854 487 397 458
Medium 9 4543 500 339 482
Long 6 8126 531 382 592
Total 25 4327 502 373 499
Table 1: For each story length bucket, we show the count (number of stories) and average length (token count) of the stories, and the average length (word count) of the summaries generated by each model for those stories.
Model Coverage Density Compression
GPT-4 72.52% 1.33 7.45
Claude 73.65% 1.26 10.86
Llama 66.09% 1.08 7.66
Table 2: We report the average coverage, density, and compression scores for the summaries generated by each model in relation to the stories.
We evaluate automatic summarization using three recent LLMs: GPT-4 (November 2023 update), Claude-2.1, and Llama-2-70B-chat (instruction-tuned for chat). GPT-4 OpenAI (2023) and Claude-2.1444https://www.anthropic.com/news/claude-2-1 have very long input contexts and can ingest an entire short story. Llama-2 Touvron et al. (2023) serves as a comparison point for smaller open-source models. Chang et al. (2023) also used these three LLMs in their work on automatic coherence evaluation in long narrative summarization.
Protecting the writers’ unpublished work is paramount when using each of these three models. We do not want their stories to be saved or trained on by OpenAI, Anthropic, or HuggingFace. We inform the writers of this risk in the consent form and complete the available forms/settings with each of these companies to request our data not be stored for long periods or used for training.
For each story, we then generate three summaries – one from each of the three models. For GPT-4 and Claude, we use one prompt (see Figure 1) as they both have a context length long enough to process a whole short story. Claude refuses to summarize two of the stories which do not meet its content restrictions [^5] For all of the Claude results, we remove the values for these two summaries to show representative numbers for the summaries it did produce. We also remove one coherence score which was inaccurate due to a bug in the annotation interface.. One of these stories is about someone being shot and the other involves sex and robbery.
Given Llama’s short context window, we use hierarchical merging of chunk-level summaries Chang et al. (2023) for stories in the medium and long length buckets. Depending on length, we chunk each story into 1-4 chunks using section or paragraph breaks. Llama summarizes each of these chunks separately and then summarizes the concatenation of these summaries (see Figure 1).
For all three models, we access them from December 2023 through January 2024, and we use simplified versions of Chang et al. (2023)’s prompts (see Figure 1). For GPT-4666gpt-4-1106-preview on https://platform.openai.com/ and Claude777claude-2.1 on https://console.anthropic.com, we use max_tokens=1000 and temperature=0.3. For Llama888meta-llama/Llama-2-70b-chat-hf on https://huggingface.co/, we use the default settings in HuggingChat [^9] https://huggingface.co/chat/.
Table 1 displays summary statistics for the generated summaries. Notice that our prompt asks models to use about 400 words and Claude undershoots but comes closest to this target while GPT-4 and Llama are 100 words over on average. See Appendix B for more prompting details and experiment costs.
We also compute the average coverage, density, and compression metrics for each model’s summaries (see Table 2). Coverage is the percent of words in the summary that come from the story, density is the average length of segments that directly match the story text, and compression is how many times shorter the summary is in relation to the story Fabbri et al. (2021a); Grusky et al. (2018). These metrics show that Llama is the least extractive and Claude is the most effective at compressing the information.
5 Human Evaluation
Figure 2: Interface screenshots showing the questions writers are asked to evaluate the summaries of their stories using a 4-point Likert scale.
We compare models using a combination of quantitative and qualitative evaluation.
5.1 Quantitative
We assess quality in terms of four attributes:
Coverage – Does the summary cover the important plot points of the story?
Faithfulness – Does the summary misrepresent details from the story or make things up?
Coherence – Is the summary coherent, fluent, and readable?
Analysis – Does the summary provide any correct analysis of some of the main takeaways or themes from the story?
Coverage, faithfulness, and coherenece are commonly evaluated in summarization (e.g., Zhang et al. (2023b)), and we add analysis which is important for capturing narrative. Each writer is shown the summaries of their own stories to evaluate as they are deeply familiar with the contents and can therefore judge aspects like faithfulness and analysis quickly and accurately. The writers rate each of these attributes on a Likert scale from 1 to 4 with some guidance on what each score means (see Figure 2, full interface shown in Appendix D).
For each story/summary pair, we also ask GPT-4 and Claude (Llama’s context window is too short) these four questions with the same wording to see how their answers compare to the opinions of the writers (see Appendix B for prompts). We additionally run BooookScore as another comparison point for LLM-based coherence evaluation Chang et al. (2023).
Once each writer has read and evaluated each summary for a story, we ask them to rank the three summaries from the three different models in order of their preference.
Lastly, we analyze the complexity of the story wording using the Flesch-Kincaid readability test Kincaid et al. (1975). This score estimates the number of years of formal education (grades 1 through 12) one might need to understand the writing easily. It is based on average word and sentence length, so while it captures something about the complexity of the wording, it is in no way a measure of the overall quality of the writing. We explore whether the reading-level of the writing affects the summary quality (see examples of writing with assigned grade levels in Figure 3).
Figure 3: Examples of openings from stories scored at different reading-levels by the Flesch-Kincaid score.
5.2 Qualitative
In addition to the ratings, the writers are also asked to provide open-ended feedback on each summary. We read all of the stories/summaries in full and assemble a list of paired text from the stories and summaries that support the writers’ lower scores and written feedback.
To examine which aspects of story content are most often missed in summaries, we ask annotators to categorize the instances of faithfulness errors from this list using elements from narrative theory for evaluating narrative understanding in children Xu et al. (2022); Paris and Paris (2003); Mandler and Johnson (1977). We define the categories for the errors as follows. An error introduces a detail that would incorrectly answer one of the narrative understanding questions in these categories:
Feeling – Emotions/reactions/internal state of characters. Answers a question like, “How did X react?” or “What was X thinking?”
Causation – Causal relationship of events. Answers a question like, “Why did Y happen?” or “What was the result of Y?”
Action – Behavior of a character. Answers a question like, “What did X do?”
Character – Identity or nature of a character. Answers a question like, “How would you describe character X?” or “Who is X?”
Setting – Story world and the time/place of events. Answers a question like, “Where/when did X happen?” or “What is the setting of the story?”
We sample 20 faithfulness errors for each model from across all its summaries. For each error, we ask three Natural Language Processing graduate students to first label whether they agree the details described constitute a faithfulness error, and if so, to categorize it (see full instructions in Appendix D). The Fleiss Kappa score is .51.51.51.51 indicating moderate inter-annotator agreement. For analysis, we use the majority decision for each error, and label it as no-consensus if multiple annotators thought it was not a faithfulness error or all three annotators gave it a different label. We also ask GPT-4 and Claude the same questions using the same prompting method as for answering the quantitative questions but with the instructions for this task (see Appendix B for prompts).
To examine how writing style affects model summarization, we use Genette’s model of narrative elements Piper et al. (2021); Genette (1980):
Narrating – Narrator’s influence on style
Story – All the events implied by the narrative
Discourse – Ordering/inclusion of explicit events
At the narrating level, we compare stories with an unreliable vs. reliable narrator. An unreliable narrator communicates something different from what the reader is meant to perceive Booth (1983). For example, one narrator states early in the story, "If I were to describe myself, I would say that I am practical. I try to be logical, but that isn’t exactly what I mean. Here’s another way to put it: there are two kinds of people in the world, me and my brother Jack." The narrator says many things without really saying anything, which gives us a clue that he may not actually be logical and practical. At the story level, we compare stories with a detailed subplot involving niche knowledge against those which focus on commonplace settings. For example, one detailed story centers on the main character sleeping with her sister’s boyfriend, but the subplot to this emotional arc involves a lot of ancient greek history. At the discourse level, we compare stories with flashbacks against those with a linear timeline. We hypothesize that reliable narrators, fewer details, and linear plots will be easier to summarize. The writer of each story verifies the labels for the story in these three categories.
6 Results
We first discuss the findings from the quantitative metrics, and then review our error analysis and the qualitative feedback from the writers.
6.1 How good are the summaries?
Model Cover. Faithful. Coheren. Analys. Avg.
GPT-4 3.48 3.12 3.52 3.40 3.38
Claude 3.17 2.65 3.41 3.26 3.12
Llama 2.40 1.92 3.08 2.76 2.54
GPT-4 56% 44% 60% 56% 54%
Claude 39% 30% 59% 43% 43%
Llama 12% 8% 32% 20% 18%
Table 3: Average scores assigned to each attribute of a summary by the writers. The first set of rows are the averaged raw scores out of 4, and the second set of rows are the percent of summaries given a perfect score of 4.
[…]
Interestingly, we find that while models have the appearance of fluency, they are given a less than perfect score for coherence 40% or more of the time as there is more to a coherent narrative summary than just fluency. In the opposite direction, the models do surprisingly well at analysis. The writers felt that 56% of the GPT-4 summaries summarized some of the themes and interpretation they had hoped to communicate as a writer, which is a challenging task even for humans (see examples of the best analysis in Figure 4).
Figure 4: Examples of some of the best analysis-focused sentences from GPT-4 and Claude summaries.
[…]
Overall, GPT-4 and Claude can produce excellent summaries but only about half the time. On the best summaries, the writers commented things like: "it did analysis that even I — the writer — hadn’t done! Very clever.", "I’m stunned at the thoughtfulness and thoroughness", and "something that I would find on a Sparknotes website".
On the worst summaries, they commented things like: "completely misses the fundamental thread", "tries to make too many analyses like a high school english class while missing the overall bigger feelings", and "makes use of…stock phrases or stand-ins rather than accurate, specific summary".
Also, for one of the stories Claude refused to summarize, the writer commented, "Art is such an important tool for processing. I don’t like that the summary wasn’t able to process a recurring vague nightmare."
Figure 5: Percent of each model’s summaries that are ranked higher than another’s in the 3-way ranking the writers are asked to do. Yellow indicates GPT-4, orange is Claude, and blue is Llama.
6.2 What mistakes do the models make?
Narrative Element
GPT-4 Claude Llama Avg.
Feeling
45% 20% 15% 27%
Causation
20% 20% 10% 17%
Action
5% 20% 25% 17%
Character
5% 5% 30% 13%
Setting
15% 15% 15% 15%
No-consensus 10% 20% 5% 12%
Table 4: Percent of faithfulness errors categorized under each narrative element for each model.
Figure 6: Examples of faithfulness errors in each of the categories assigned by the annotators.
In Figures 6 and 7, we show examples of issues models have in each attribute – coverage, faithfulness, coherence, and analysis. In general, we find that while writers rank GPT-4 and Claude summaries somewhat equally, Claude summaries often leave out important details that lower their coverage score and create issues with faithfulness as identified by the numerical scores. For example, while the GPT-4 coverage error in Figure 7 oversimplifies a major plot point in the story, the Claude error removes an entire thematic subplot and creates a faithfulness error that the mother is still ill instead of recovered.
By categorizing the faithfulness errors in Table 4, we find that most faithfulness errors are in Feeling when averaged across the models. These errors require interpretation of character reactions and internal states, so this result demonstrates models still struggle with reading people. However, the error distribution for each model varies. For example, Llama has a high percentage of Character errors. This is likely an artifact of the chunk-then-summarize strategy as sometimes characters are conflated or split in two as the model fails to track their details across different chunk summaries (see Character example in Figure 6). Chang et al. (2023) also found a similar type of error (entity omission) to be most common with chunk-then-summarize approaches.
It is interesting to note that some faithfulness errors come from the models being overly normative as evidenced by the Claude Feeling error in Figure 6. In this case, Claude assumes a heteronormative monogamous interpretation of the interaction instead of what is written. One writer observed this phenomenon in another summary, commenting that the summary had "a few extrapolations… which presume things about the narrative based on a hypernormative interpretive framework".
Lastly, some of the writers commented that the models would copy direct and unique wording from the stories without placing it in quotations, which bordered on plagiarism. We flag this for future work.
Figure 7: Examples of coverage, coherence, and analysis errors the models make.
6.3 Do some aspects of writing style affect summary quality?
6.3.1 Length
Figure 8: Averaged writer scores for summaries for stories within a given length bucket for each model. The bar saturation indicates the length bucket.
In Figure 8, we compare the average scores across different story length buckets. The stories in the short bucket are short enough that Llama can fit the entire story in its input context. Looking at this bucket, the scores are better for Llama than the longer buckets but still not better than the other two models. For Claude and GPT-4, it seems the long-context models can summarize short and long stories equally well (up to 10,000 tokens). For these two models, the summaries of the longest stories have the highest average scores by a slim margin, but this may be because there are fewer stories in this bucket (only 6).
6.3.2 Reading-Level
We plot the reading-level (estimated years of education needed to understand the wording easily) against the writer-assigned scores averaged across the four attributes for each summary. We find no significant correlation between story reading-level and writer-assigned summary scores for any of the models (GPT-4: $r$11, $p$61; Claude: $r$33, $p$12, Llama: $r$23, $p$27). The average Flesch-Kincaid grade-level for the group of stories is 6th grade. As you can see in the 2nd grade example in Figure 3 though, which features someone being shot, this score captures an aspect of the wording but does not consider age-appropriate content or conceptual complexity.
6.3.3 Narrating, Story, and Discourse
Style Element # GPT4 Claude Llama
Narrating: reliable 15 3.45 3.29 2.60
unreliable 10 3.28 2.89 2.45
Story: detail-- 16 3.42 2.98 2.59
detail$+$ 9 3.31 3.36 2.44
Discourse: linear 18 3.29 3.05 2.56
nonlinear 7 3.61 3.32 2.50
Table 5: Average scores for summaries of stories with different style elements. Each score is averaged across the four attributes for a summary.
In Table 5, we see that stories with unreliable narrators are harder for all three models to summarize. All three models have lower average scores in this bucket. As an example of an error due to an unreliable narrator, the same unreliable narrator who is quoted in Section 5.2 further describes himself as wanting a "small, normal life", and then describes his new love interest:"I met a girl… I think she’s just like me. She loves traveling… It’s pretty cool." GPT-4 mistakenly characterizes this interaction as: "The protagonist meets [a girl]… with whom he shares… a desire for a simple life." even though there is no evidence she likes a simple life (or even that we should believe the narrator that he does), other than that the protagonist has described her as "like [him]".
For story and discourse, the models have mixed performance across the different buckets. Llama does have lower scores in the buckets we hypothesized might be more challenging. Overall, it seems that unreliable narrators are a challenge for LLMs, whereas level of detail and flashbacks depend on the case. One writer commented directly on the issue of an unreliable narrator by observing, "there’s a second layer here that the summary misses somewhat—the narrator is not really to be trusted".
6.3.4 Can an LLM replace the writers in evaluating the summaries?
Figure 9: Confusion between model-assigned scores/labels and human-assigned scores/labels for GPT-4 and Claude.
Attribute GPT-4 Claude
Summary Rating: Coverage .46* .18
Faithfulness .37* .05
Coherence .18 .02
Analysis .21* .02
Faithfulness Error Labels .65 .53
Table 6: Agreement between model-assigned scores/labels and writer-assigned scores/labels. For the summary ratings which are Likert scores, we report Pearson’s $r$ (* indicates p-value <.05). For the faithfulness errors which are categorical, we report Cohen’s Kappa.
In Figure 9, we compare the GPT-4 and Claude scores to the writer scores from Table 3, and we see considerable confusion across the ratings (first row of plots). Claude mostly only rates summaries as 3 or 4, causing overestimation of ratings (3.67 average score vs. 3.01 from the writers).
In terms of attributes, both models rate the coherence of the summaries much higher than the writers do (average scores: GPT-4 - 3.84, Claude - 3.82, Writers - 3.33). GPT-4 largely underestimates the coverage score for all three models (average score 2.68 vs. 3.01 from the writers). Many of the writers commented that the GPT-4 summaries were too detailed, so the model may struggle with identifying what is most important to cover. Claude overestimates performance on faithfulness and analysis. It gives 85% of summaries a 4 on faithfulness on average and 85% on analysis as well, relative to averages 27% and 39% respectively from the writers. GPT-4 also overestimates faithfulness performance just for itself (64% scores of 4 vs. 44% from the writers).
We also show the Pearson’s $r$ correlation between the model scores and the writers’ scores in Table 6 across attributes. We see no correlation for Claude scores and some statistically significant but weak correlation for GPT-4 scores, other than for coverage which has moderate positive correlation. We also find no significant correlation between BooookScore coherence scores and the writer coherence scores ($r$:.11, $p$:.18). The average BooookScore scores are almost the same across all three models (GPT-4: .95, Claude: .93, Llama: .93), which does not capture the significant difference in coherence that the writers judged in Llama in particular. Overall, these results support prior work Chakrabarty et al. (2023a), which shows that LLMs are not yet reliable evaluators for skilled writing tasks.
However, we do find that GPT-4 and Claude have moderate to substantial agreement with our human labelers for categorizing the faithfulness errors (see Table 6 and the second row in Figure 9). Importantly, they are not actually detecting the faithfulness errors themselves, but performing this error categorization could be helpful for future error analysis in studies like this.
7 Discussion
Our results show that models are capable of thorough and interesting analysis of the themes present in unseen stories. When models are asked to summarize stories that may have analysis available online, it isn’t clear if the models are simply regurgitating analysis from the training data. Therefore, it is important to examine their capabilities on original and challenging content. Additionally, while we expected it might be challenging for LLMs to identify salient information within long stories, we find that long-context models demonstrate as good understanding of longer stories as shorter ones.
The models also have notable weaknesses. We demonstrate that even the best models are making significant errors across all summary attributes on about half of summaries. Many of the faithfulness errors come from misinterpreting challenging subtext and particularly character feelings and reactions. This will be an important area for future study as it demonstrates language models still struggle with aspects of theory of mind [^10] Theory of mind is the ”ability to understand and keep track of the mental states of others” Xu et al. (2024)., one of the most challenging being understanding an unreliable narrator. Additionally, our analysis of LLM-based evaluation shows that while LLMs can assist with evaluation, they should not replace the expert human judgments.
Working with writers was an efficient and mutually enjoyable method of ensuring we were not evaluating on any training data, and we were getting an informed evaluation of characteristics like faithfulness and analysis. Writers were able to complete summary ratings in a matter of minutes, whereas someone unfamiliar with the story would have taken much longer. For example, it took us, the paper authors, about 60-90 minutes to read and review the summaries for just one story, whereas it took the writers abot 5-10 minutes to do the same. Writers also left positive feedback like, "I’m glad to have read this [summary]… It shows some [weaknesses] of my story… some minor characters are more flat than I want."
Limitations. We provide a first look at how LLMs summarize unseen short stories within a reasonable cost and timeframe, but the work is limited in its number of stories, ratings, and prompting techniques. A follow-up study could expand these areas, but it is challenging given that an individual author has a limited number of high-quality stories that are complete but unpublished, there are a limited number of writers willing to participate in a study involving LLMs and corporate APIs, and compensation is expensive. As a result of these challenges, our numbers of stories and writers are similar to other studies which have used experienced creative writers as shown in Section 2.
Ethics. We follow protocol approved by our IRB for this study. In line with the TACL code of ethics, we protect writers’ work and identities by saving data on secure servers attached only to anonymous IDs, approving any published excerpts from their work with them, and requesting our prompting data not be used for future model training. We also support the code of ethics by involving practitioners from the field influenced by our study in our work. One of the authors, Melanie subbiah, has an equity interest in OpenAI.
8 Conclusion
We work with writers to provide unpublished short stories and evaluate the quality of LLM-generated summaries of these stories. We present a quantitative and qualitative holistic evaluation of summary quality grounded in narrative theory that is based on data LLMs did not train on. In doing this, we identify that LLMs can demonstrate understanding of long narrative and thematic analysis, but the remaining challenges lie in reliable interpretation of subtext - specifically feeling and narrative voice. Our methodology sets an important example of how we can collaborate with domain experts to reach beyond the paradigm of evaluating LLMs with LLMs on data they may have been trained on.
Appendix A Information Sheet for Informed Consent
Purpose of research study: We are collecting judgments from writers of how well automatic summarization models can summarize the writers’ own short stories to study the level of understanding these models demonstrate of this type of writing.
Benefits: This study may not benefit you directly, although it may be interesting to see an AI model’s takeaways from your story. The study could, however, benefit society through identifying strengths/limitations of AI models when processing nuanced long text documents.
Data protection: We understand it is very important to protect your unpublished short stories. We will store your stories on password-protected Columbia University computers for the duration of this study and then delete them after completion of the study. We may like to include short excerpts (<1 paragraph) from your short stories and evaluation responses in an eventual publication as examples, but we will not do so without getting your permission at that time. Other than this, we will not publish your stories in any way. We do have to run your stories through the OpenAI, Anthropic, and HuggingFace company APIs to access some of the AI models. All three companies state that they will not train on any data. The data may be retained at the company for 30-90 days to ensure consistency in their service but this data is not used or released by the company.
Risks: Since, we have to run your unpublished short story through the OpenAI, Anthropic, and HuggingFace company APIs, there is a small risk involved with trusting these companies are doing what they say they’re doing and preventing data leaks. Anthropic is additionally HIPPAA certified for safe data handling, and one of the authors, Melanie Subbiah, has run her own unpublished short stories through these APIs.
Voluntary Participation: You may stop participating at any time without penalty by emailing {PI} ({PI-email}). At that time, we will delete any short stories you may have shared with us. You will only be compensated for stories for which you complete the full evaluation process. We may end your participation if you are not following the instructions.
Compensation: You will be compensated \$50 for each short story/evaluation you complete in the form of a prepaid VISA card mailed to you.
Confidentiality: No identifying information will be published about you. We record whether you have published a short story before and/or completed a writing program to give a general sense of our study participants. Your mailing address will be kept for the duration of the study so that we can mail your payment card to you, and then promptly deleted. All of your work will be submitted in relation to an anonymous ID so we will not be able to connect your answers to you. The Columbia University Human Research Protection Office may obtain access to the data collected for this study.
COI Disclosure: Please be aware that Melanie Subbiah (an investigator on this study) has a financial interest in OpenAI, a company which makes some of the automatic summarization models which will be evaluated in this study.
Questions/concerns: You may email questions to the principal investigator, {PI}: {PI-email}. If you feel you have been treated unfairly you may contact the Columbia University Institutional Review Board and reference Protocol AAAU8875. Please respond to this email by copying the below consent paragraph to indicate that you understand the information in this consent form. You have not waived any legal rights you otherwise would have as a participant in a research study. If you have any questions about your rights or responsibilities as a research participant, please contact the Columbia University HRPO office at: Phone {phone-number}; Email {email}. We recommend either printing this information sheet or taking note of this contact information before beginning the study in case of questions.
“I have read the above purpose of the study, and understand my role in participating in the research. I volunteer to take part in this research. I have had a chance to ask questions. If I have questions later, about the research, I can ask the investigator listed above. I understand that I may refuse to participate or withdraw from participation at any time. The investigator may withdraw me at his/her professional discretion. I certify that I am 18 years of age or older and freely give my consent to participate in this study.”
Appendix B Model Prompts and Costs
B.1 GPT-4
Cost (experiments, summary generation, and evaluation): \$58
Summary Generation:
System prompt
You are an expert summary-writer. Summarize the provided passage in several paragraphs using only information from the passage provided.
User prompt
[…]
Summary Evaluation:
System prompt
You are a skilled writer and editor. Evaluate the quality of a summary of a short story by answering the questions provided. Select from the four options indicated and choose whichever fits best. Be careful to evaluate the summary in relation to the short story provided.
User prompts (questions are asked one by one and model answers are kept in the messages)
[…]
Faithfulness Error Labeling:
System prompt
{instructions loaded from same file as for human annotators}
User prompts (questions are asked one by one and model answers are kept in the messages)
[…]
B.2 Claude
Cost (on Claude limited free plan): \$0
from anthropic import HUMAN_PROMPT, AI_PROMPT
Summary Generation:
Prompt
{HUMAN_PROMPT} {gpt4_user_prompt}{AI_PROMPT}
Summary Evaluation:
Prompt (questions asked one at a time in this format)
[…]
{AI_PROMPT}
Faithfulness Error Labeling:
Prompt (questions asked one at a time in this format)
[…]
{AI_PROMPT}
B.3 Llama
Cost (on HuggingChat): \$0
Summary Generation:
System prompt
You are an expert summary-writer. Summarize the provided story in several paragraphs using only information from the story provided.
Story-level (if the story is short enough) prompt:
Write a coherent, chronological, and detailed summary for this story in several paragraphs (about 400 words). Briefly introduce key entities like characters or settings when they are mentioned in the summary, and include some analysis of the story.
Chunk-level prompt:
Write a coherent, chronological, and detailed summary for this part of a story in several paragraphs (about 400 words). Briefly introduce key entities like characters or settings when they are mentioned in the summary, and include some analysis of the story.
Concatenated-summaries prompt:
Compress these summaries of several parts of a story into a couple paragraphs (about 400 words) that are a coherent, chronological, and detailed summary of the whole story. Briefly introduce key entities like characters or settings when they are mentioned in the final summary, and include some analysis of the story.
Appendix C Short Story Uploader and Evaluation Interfaces
See Figure 10.
Cost: \$1,250 to compensate writers for their work
Figure 10: Screenshots of the Streamlit apps writers use to anonymously submit stories and evaluate summaries.
Appendix D Faithfulness Error Annotation Instructions
See Figure 11.
Figure 11: Annotator instructions and an example error for annotation of faithfulness error labels.